The center of a circle is the origin of a rectangular coordinate plane. If (-4,0), (0,4) and (4,0) are three points on the circumference of the circle, what is the probability that a randomly picked points inside the circle would fall inside the triangle formed by those three points?
a. 1/2
b. 1/3
c. 1/pi
d. 2/pi

Answer by Boreal(15235)


The area of the circle is pi*r^2=16 pi
The area of the triangle is 1/2 bh, where the base is 2r=8 and h=4. That product is 16.
16/16 pi=1/pi